OpenStudy (anonymous):

Can someone give me the answer please. The table below shows two equations: Equation 1 |4x − 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? A)Equation 1 and equation 2 have no solutions. B)Equation 1 has no solution, and equation 2 has solutions x = −4, 1. C)The solutions to equation 1 are x = 3, −1.5, and equation 2 has no solution. D)The solutions to equation 1 are x = 3, −1.5, and equation 2 has solutions x = −4, 1.

OpenStudy (lynfran):

solution 1: \[4x-3-5=4\]\[4x-8=4\]\[4x=12\]\[x=3\] solution 2: \[4x+6+8=3\]\[4x=-11\]\[x=-\frac{ 11}{ 4 }\]

OpenStudy (anonymous):

so the answer would be D? @LynFran

OpenStudy (lynfran):

For equation 1; when you substitute x=3 you get 4, however substituting x=-1.5 you get -14 and for equation 2; \[x=-2\frac{ 3 }{ 4 }\] not -4 or 1

OpenStudy (lynfran):

the equationshave solution which are x=3 for 1st equation and \[x=-2\frac{ 3 }{ 4 }\] for the 2nd equation